A body starts from rest and has constant acceleration of 2m/s 2. Find its velocity at t=2 seconds
1
Expert's answer
2020-05-15T08:41:08-0400

Velocity of the object, moving with constant acceleration, at time "t" is "v(t) = v_0 + a t", where "v_0" is initial speed and "a" is acceleration.

Hence, "v(2) = 0 \frac{m}{s} + 2 \frac{m}{s^2} \cdot 2 s = 4 \frac{m}{s}".